OpenStudy (anonymous):

Solve for x: 3/4 X + 5/8 = 4x

OpenStudy (akashdeepdeb):

\[\frac{3x}{4} + \frac{5}{8} = 4x\] \[\frac{6x + 5}{8} = 4x\] \[6x + 5 = 8 * 4x\] \[6x + 5 = 32x\] \[x = \frac{5}{26}\]

OpenStudy (anonymous):